OpenStudy (anonymous):

paula works between 10 and 15 hours a mouth . if she charges $4.50 an hour , what is a reasonable amount of money she can earn in 3 mouth ? and graph

OpenStudy (anonymous):

Any multiple choices?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

$45-67.60

OpenStudy (anonymous):

that the answer

OpenStudy (anonymous):

somewhere between \[3\times 10\times 4.50\] and \[3\times 15\times 4.50\]

hero (hero):

Chris, look like you've got plenty help here

OpenStudy (anonymous):

say $135 for the lower estimate and $202.50 for the upper

OpenStudy (anonymous):

can you help me graph it

OpenStudy (mertsj):